please help me, i need help with this: 1) {8-5}2+5^2 2) {10-6}+2^2 3) 7{12-2{3+2}
Found 2 solutions by ikleyn, addingup:
Answer by ikleyn(52788)

.
1)  {8-5}2 + 5^2 = ?

    (8-5)*2 = 3*2 = ?  

    5^2 = 5*5 = 25. 

    Complete on your own from this point.



2)  (10-6) + 2^2 = ?

     10 - 6 = ?

     2^2 = 2*2 = ?

    Complete on your own from this point.


3)  is written incorrectly.


Answer by addingup(3677)

1)
{8-5}2+5^2 use the Distributive Property of Multiplication to distribute the 2 over the first parenthesis, square the 5, and add everything up:
= 16-10+25 = 31
------------
2)
To solve, remember PEMDAS (BODMAS if you are in Canada, the U.K., Australia, etc).
PEMDAS: Parentheses, Exponents, Multiplication and Division, and Addition and Subtraction
BODMAS: Brackets, Orders, Division and Multiplication, and Addition and Subtraction.
{10-6}+2^2 Solve inside parenthesis, the exponent, and add:
= 4+4 = 8
---------------------
3)
Watch your signs! (+, -)
7(12-2(3+2))
= 7(12-6-4)
= 7(2)
= 14